OpenStudy (anonymous):

Ask a Question..♥♥♥☻ Expand 27x3+125y3. ☺(3x+5y)(9x2-15xy+25y2) ☺(3x-5y)(9x2-15xy+25y2) ☺(3x+5y)(9x2+15xy+25y2) Which One is Correct

Parth (parthkohli):

Are you expanding or factoring?

OpenStudy (anonymous):

Asking..

hartnn (hartnn):

Expand 27x3+125y3. 2 7 x 3 + 1 2 5 y 3 .

Parth (parthkohli):

Oh, and\[\rm a^3 + b^3 = (a +b)(a^2 - ab + b^2)\]

OpenStudy (anonymous):

I think expanding means the same as factoring here.

OpenStudy (anonymous):

Good to note here that 27x^3+125y^3 = 3^3x^3 + 5^3y^3 = (3x)^3 + (5y)^3

Parth (parthkohli):

And then use that silly thing I wrote in my last reply.